OpenStudy (anonymous):

A dog eats 5 cups of dry dog food in 2 1/2 days. How many cups of dry dog food does the dog eat in 1 day? cups

OpenStudy (anonymous):

Divide 2 1/2 / 5.

OpenStudy (anonymous):

what? how?

OpenStudy (anonymous):

2 1/2 is the same as 2.5.

OpenStudy (anonymous):

So just divide 2.5 / 5

OpenStudy (anonymous):

it is 0.5

OpenStudy (texaschic101):

5/2.5 = x/1 --- 5 cups in 2.5 days = x cups in 1 day cross multiply 2.5x = 5 x = 5/2.5 x = 2 so in 1 day, he eats 2 cups

OpenStudy (anonymous):

@texaschic101 Umm... That's wrong.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Wait, I think I'm wrong. LOL

OpenStudy (anonymous):

Oh, you have to divide 5 / 2.5, not 2.5 / 5... (FaceDesk)

OpenStudy (anonymous):

My mistake, @texaschic101 is correct.

OpenStudy (anonymous):

so the answer is 2 cups?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I need more help please

OpenStudy (anonymous):

Dora can run 3/5 mi in 3/20 h. How far can she run in 1 h? mi

OpenStudy (anonymous):

Please ask your questions in separate threads.

OpenStudy (anonymous):

ok

OpenStudy (uri):

Easy,If a dog eats 5 cups of dry food in 2.5 days,He will eat less cups of dry food in 1 day,so direct proportion,Cross multiply. For the first one @cjackson2002